OpenStudy (anonymous):

What is the value of 1/2^-2x^-3y^5 for x=2 and y=-4? SIMPLIER THAN IT LOOKS!!!

OpenStudy (anonymous):

a. 16 b. -4 c. -1/32 d. -16

OpenStudy (anonymous):

@ganeshie8 help? I been trying to work this thing out multiple ways and every single answer I get is freaking not on the answer choices and urgh I am really good at this but I juts can't figure this one out. help? please

OpenStudy (anonymous):

It's a negative and zero exponent question, really, it's soo simple dam

OpenStudy (anonymous):

thanks for looking at my question, by the way!!^^

OpenStudy (anonymous):

want me to show you what I got ?

OpenStudy (anonymous):

i got -1024/32 so I was thinking it was c casue that was the closest

OpenStudy (superdavesuper):

plz check ur expression "1/2^-2x^-3y^5 for x=2 and y=-4" two terms 1/2^ and 2x^ seem to be missing their exponents

OpenStudy (anonymous):

oh, im sorry i'll correct it right away! let me see..

OpenStudy (anonymous):

no no, that's it. it's 1/2^-2*x^-3* y^5

OpenStudy (anonymous):

it's a negative exponent problem where you inverse the negative exponent first to make it positive and then simplify it.

OpenStudy (superdavesuper):

so is it ((((1/2)^(-2x))^-3y^5)? that would be crazy complicated though...

ganeshie8 (ganeshie8):

if u have the pic, please attach... it wil be helpful :)

OpenStudy (anonymous):

yes, give me a second! ^^

OpenStudy (anonymous):

this is the problem!

OpenStudy (anonymous):

I'm sorry if that's not helpful, umm, maybe if you show you what i mean by inverse of the negative exponent to turn them to positive fractions

OpenStudy (anonymous):

give me one sec, k, and thank you both for trying to help!!!^^

OpenStudy (superdavesuper):

no, its much better because what u typed before doesnt make any sense.

OpenStudy (superdavesuper):

now stop for a moment n try to do it my way ok?

OpenStudy (anonymous):

oh, I'm sorry!!!

OpenStudy (anonymous):

ok!

OpenStudy (superdavesuper):

ignore the 1/ part and just look at 2^-2 * x^-3 * y^5 okay? put x=2 n y=-4 or y=-2^2 in there plz

OpenStudy (anonymous):

I did, want me to solve it and give you the answer?

OpenStudy (anonymous):

simplify it and report back?

OpenStudy (superdavesuper):

no, just show ur steps here plz i will do the 1st part: 2^-2 * 2^-3 * (-2^2)^5 plz continue from there

OpenStudy (anonymous):

ok! welllll, since the whole lesson is about negative exponents and zero exponents, I have to inverse the negative exponent to a positive so I'll turn the negative to positive like this: 2^-2 = 1/2^2 and 2^-3 = 1/2^3 is that right though?

OpenStudy (anonymous):

1/4 * 1/8 * (-2^2)^5

OpenStudy (superdavesuper):

u will be drawing urself crazy if u do that though....it would be much simpler if u just follow the steps above and stay....oh well....ok plz go on....

OpenStudy (anonymous):

ok! 1/4 * 1/8 * 24 (i got 24 from the 2^5 then simplified it and got 24) = 24/32 yes. any thoughts on how else this problem could be solved cause this is obviously not one of the answers and wow I feel really stupid right now cause I can't figure out his one problem and yeah.

OpenStudy (superdavesuper):

first 2^5 is not 24 and also a - sign seems to be missing

OpenStudy (superdavesuper):

2^5 = 2*2*2*2*2

OpenStudy (anonymous):

sorry, it's 32, i work it out on paper!

OpenStudy (anonymous):

just now i check it on calculate, sorry about that!

OpenStudy (anonymous):

Wouldn't it be -1/32?

OpenStudy (superdavesuper):

scroll up n see what exactly is the expression u are trying to calculate plz.... it should be something like (-2^2)^5...i think u are rushing through ur cal....n missed some terms n made some mistakes. there is a valid ans in ur question there :)

OpenStudy (anonymous):

Did you two already move up 2^-2 and x^-3? because you cant have a negative power right?

OpenStudy (anonymous):

you divide -1024 and 32 by 32

OpenStudy (anonymous):

OMG!!! THAT'S WHAT I GOT AT FIRST AND WOW I DIDN'T DIVIDE IT!!!

OpenStudy (anonymous):

Hey that's good (: so the answer would by the third option -1/32

OpenStudy (anonymous):

Resolved?

OpenStudy (anonymous):

THANK YOU ALL WOWOW!! IM SOO HAPPY!

OpenStudy (anonymous):

I COULD CRY, im soo tired thank you for helping guys!!!!!

OpenStudy (anonymous):

i had it! wow

OpenStudy (anonymous):

Yay (: just remember to simplify it lol

OpenStudy (superdavesuper):

good job @Duizhang✨ and @speckleleckle

OpenStudy (anonymous):

kinda had it, anyway thank you!!!!